2) The pH of coffee is approximately 5.0. How many times greater is the [H] in coffee than in neutral water? Ans: 3) Calculate the H ion concentration in a 8.8 x 10-4M Ca(OH), solution. Ans: 4) Consider the weak acid CH COOH (acetic acid). If a 0.048 M CH:COOH solution is 5.2%) ionized, determine the [HO] concentration at equilibrium.
Answer #1

Ans 2 :

[H+] = 10-pH

For coffee , where the pH is 5 , the hydrogen ion concentration will be :

[H+] = 10-5 M

For water , the pH is 7 , so the hydrogen ion concentration will be :

[H+] = 10-7 M

Diving the H+ concentration of coffee by water :

= (10-5 M) / (10-7 M)

= 100

So 100 times greater is the [H+] in coffee than in neutral water.